OpenStudy (anonymous):

If ∠m ≅ ∠p, m∠m = 9x - 2, and m∠p = 4x + 18, how many degrees are in ∠p?

OpenStudy (cggurumanjunath):

figure?

OpenStudy (anonymous):

so they are equal

OpenStudy (anonymous):

9x-2=4x+18

OpenStudy (anonymous):

9x-4x=18+2 5x=20 x=4

OpenStudy (anonymous):

<p=4*4+18=16+18=34 degree

OpenStudy (anonymous):

thank youuu

OpenStudy (anonymous):

Identify the hypothesis of the following statement: The plants will die if it does not rain soon